find the point on the curve y= (x-1)^1/2 that is closest to point (2,0)
Answer by Alan3354(69443)

find the point on the curve y= (x-1)^1/2 that is closest to point (2,0)
--------
Simplify:
find the point on the curve y= x^1/2 that is closest to point (1,0)
-----
Points on the curve are (x,sqrt(x))
---
d^2 = diffy^2 + diffx^2 = (sqrt(x) - 0)^2 + (x-1)^2 = 2x^2 - 2x + 1
------
d = sqrt(2x^2 - 2x + 1)
dd/dx = (1/2)*(2x^2 - 2x + 1)^(-1/2)*(4x-2) = 0
2x-1 = 0
x = 1/2
==============
Point (1/2,sqrt(1/2))
Move it all 1 to the right
--> point is (3/2,sqrt(1/2)
-----------
Distance = sqrt(3)/2, btw